OpenStudy (anonymous):

8x - 1 = -4x + 23 x =

OpenStudy (uri):

8x - 1 = -4x + 23 8x+4x =23 +1 12x =24 x =24/12 x =2 ans.

OpenStudy (anonymous):

Cool @uri xD